The wasy^{∗}symbol fonts for use with L^{A}TEX 2ε

Axel Kielhorn

94/05/28

1 Introduction

This file defines the package wasysym which makes some additional characters available that come from the wasy fonts (Waldis symbol fonts). These fonts are not automatically included in the NFSS2/L^{A}TEX 2ε since they take up important space and aren’t necessary if one makes use of the packages amsfonts or amssymb.

Warning: This style uses verion 2 of the wasy-fonts. It is not 100 % compatible to the old version 1 from 1989. We have provided no compatibility mode for the old fonts! If some characters come out wrong you have to upgrade.

The commands defined by the wasysym package are:

\Join ⨝ \Box □ \Diamond ◇ \leadsto ↝ \sqsubset ⊏
\sqsupset ⊐ \lhd ⊲ \unlhd ⊴ \LHD ◄ \rhd ⊳ \unrhd ⊵
\RHD ► \apprle ≲ \apprge ≳ \wasypropto ∝ \invneg ⌐
\ocircle ◯ \logof ⍟ \varint ∫ \iint ∬ \iiint ∭
\varoint ∮ \oiint ∯

The following commands are available in text-mode and scale according to the textfont size:

General symbol:

\male ♂ \female ♀ \currency ¤ \phone ☎ \recorder ⌕
\clock 🕒 \lightning ↯ \pointer ⇨ \RIGHTarrow ► \LEFTarrow ◄
\UParrow ▲ \DOWNarrow ▼ \diameter ⌀ \invdiameter ⍉
\varangle ∢ \wasylozenge ⌑ \kreuz ✠ \smiley ☺ \frownie ☹
\blacksmiley ☻ \sun ☼ \checked ✓ \bell 🔔 \ataribox
\cent   \permil   \brokenvert ¦ \wasytherefore ∴ \Bowtie ⨝
\agemO ℧

Electrical and physical symbols:

\AC ∼ \HF ∼ ∼ \VHF ≋ \photon ∼∼∼∼ \gluon

Polygons and Stars:

\Square □ \XBox ⊠ \CheckedBox □ ✓ \hexagon ⎔ \varhexagon ⬡
\pentagon ⬠ \octagon      \hexstar ✶ \varhexstar ✶
\davidsstar ✡

∗wasy-font by Roland Waldi, Universität Karlsruhe, Germany, Copyright 1989, Version 2 Copyright 1992

Music notes:

\eighthnote ♪ \quarternote ♩ \halfnote 𝅗𝅥 \fullnote 𝅝
\twonotes ♫

Various circles:

\Circle ◯ \CIRCLE ● \Leftcircle      \LEFTCIRCLE ◖
\Rightcircle    \RIGHTCIRCLE ◗ \LEFTcircle ◖ ◯ \RIGHTcircle ◗ ◯
\leftturn ↺ \rightturn ↻

Phonetic signs:

\thorn   \Thorn   \dh   \Dh D \openo   \inve

Some symbols used in astonomy:

\vernal ♈ \ascnode ☊ \descnode ☋ \fullmoon ◯ \newmoon ●
\leftmoon ☾ \rightmoon ☽ \astrosun ⊙ \mercury ☿
\venus ♀ \earth ♁ \mars ♂ \jupiter   \saturn   \uranus
\neptune   \pluto

Astrological symbols and the zodiacal symbols:

\aries ♈ \taurus    \gemini    \cancer  \leo ☊ \virgo
\libra   \scorpio   \sagittarius   \capricornus   \aquarius
\pisces   \conjunction   \opposition

And the APL-symbols, whoever needs them:

\APLstar ★ \APLlog ⍟ \APLbox ⎕ \APLup △ \APLdown ▽
\APLinput ⍞ \APLcomment ⍝ \APLinv ÷ ⎕ \APLuparrowbox ⍐
\APLdownarrowbox ⍗ \APLleftarrowbox ⍇ \APLrightarrowbox ⍈
\notbackslash \ − \notslash / − \APLnot ∼\APLcirc ∘ \APLvert | \APLminus −